SA13. A basketball player throws a desperation shot at the buzzer from the backcourt 1.0 second before the game ended with the score tied. If the ball was released at a velocity of 12.0 m/s at an angle of 38O and if the release was made at the same level as the basket from a distance of 14.2 m, did he score? Why?


1
Expert's answer
2021-11-29T11:47:17-0500

He made a throw et the distance "l=\frac{v^2\sin 2\alpha}g=14.1~m,"

but the radius of the basket is 23 cm,

so the distance is "d'=d-r=14.2-0.23=13.97<14.1=l,"

he did a score.